MCC did the beam line test with the following:
Septum_Current Corrector_Current 77.5 1.0 232.5 2.2 310 3.4 387.5 5.7 465 8.1 500 9.9
Since the maximum current for the corrector is 10 A, 500 A will be the maximum current for the septum until the permanent magnet is installed (tomorrow). After the permanent magnet is installed, the corrector will need to be re-calibrated (time TBD).
The septum current 500 A had been tested a few times, we saw the spot at the same position at the beam dump.
We also tried to see a spot using the BeO target while both Septum and corrector are zeros. We did see a spot on the target. With the septum at 500 A and the corrector at 9.9 A, the spot on BeO was near the edge of the BeO window.